### v2

# Variations on quantum de Finetti theorems and operator valued Martin boundaries: a Choquet–Deny approach

Abstract.
We revisit the quantum de Finetti theorem. We state and prove a couple of variants thereof. In parallel, we introduce an operator version of the Martin boundary on quantum groups and prove generalizations of Biane’s theorem. Our proof of the de Finetti theorem is new in the sense that it is based on an analogy with the theory of operator valued Martin boundary that we introduce.
1. Introduction
The quantum de Finetti theorem is a fundamental theorem in quantum information theory. It states that the collection of separable states corresponds precisely to the collection of infinitely extendable symmetric states. It was developed independently in operator algebras [12] and in quantum information theory [7]. This remains one of the most powerful tools to detect entanglement: a state is entangled if and only if it fails to be $k$-extendable for $k$ large enough. In general, proofs of this phenomenon rely heavily on the representation theory of the symmetric group.
In this paper, we propose a new approach to this theorem, inspired by the theory of random walks on commutative groups, and, more precisely, on the notion of the Martin boundary and the Choquet-Deny theory. In these theories, the harmonic elements, which are considered, are scalar-valued functions because studying more general functions do not enrich the theory. In their non-commutative generalizations, Biane and Izumi did not consider operator-valued boundaries. In this paper, we study operator-valued boundaries and observe that it is closely related to the problem of identifying separable states.
After a short section on notations, we provide in Section 3 a new proof of the generalized quantum de Finetti theorem, inspired by harmonic analysis. It mimics Biane’s proof of the non-commutative Choquet-Deny theorem.
In Section 4, which seems a priori unrelated to the preceding section, we study the matrix-valued Martin boundary on the dual of a compact (quantum) group. This extends the previous workof Biane [1]. Then we observe a phenomenon comparable to the one in Section 3: the matrix-valued Martin boundary is the separable version of its non-matrix-valued counterpart.
In Section 5, we explore further the relation between sections 3 and 4 by showing an inclusion between the matrix-valued non-commutative Martin boundary and the collection of extendable matrix-valued states. This inclusion is based on a matrix-valued extension of a non-commutative Martingale convergence theorem considered by Biane and Izumi.
The last section delves into the problem of understanding the notion of exponentials better. We extend some results of Biane [2] to more groups, and in the matrix-valued setup. This allows us to identify cases when the injection described in part 3 is a bijection.
B. C. is supported by JSPS Grant-in-Aid Scientific Research (B) no. 21H00987, and Challenging Research (Exploratory) no. 20K20882 and 23K17299.
T. G. is supported in part by a grant from NSERC, Canada.
R. S. was funded by JSPS Research Fellowship for Young Scientists PD (KAKENHI Grand Number 22J00573).
2. Notations
In this section, we summarize the notations and conventions that we use throughout this paper. Let $\mathbb{M}_{n}(\mathbb{C})$ denote the *-algebra of $n\times n$ matrices over $\mathbb{C}$ and $\mathrm{Tr}_{n}$ the (non-normalized) trace on $\mathbb{M}_{n}(\mathbb{C})$. We denote by $\mathbb{M}_{n}(\mathbb{C})_{+}$ the set of all positive semi-definite matrices. In the paper, elements in $\mathbb{M}_{n}(\mathbb{C})_{+}$ are simply called positive. More generally, let $B(\mathcal{H})$ denote the *-algebra of bounded linear operators on a complex Hilbert space $\mathcal{H}$. Here, its multiplication is given as compositions of linear operators and its *-operation is given by taking adjoint operators. Let $\mathcal{A}\subseteq B(\mathcal{H})$ be a unital $C^{*}$-algebra, i.e., $\mathcal{A}$ is a unital *-subalgebra of $B(\mathcal{H})$ and closed with respect to the topology induced by the operator norm. We denote by $\mathcal{A}_{+}$ the set of all positive elements in $\mathcal{A}$, that is, by definition, $x\in\mathcal{A}_{+}$ if and only if $x$ is self-adjoint and the spectrum of $x$ is contained in $\mathbb{R}_{\geq 0}$. For any pair $x,y\in\mathcal{A}_{+}$ we write $x\leq y$ if $y-x\in\mathcal{A}_{+}$. We remark that $\mathcal{A}_{+}$ is a convex cone.
Let $\mathcal{A}^{*}$ denote the set of all (continuous) linear functionals on $\mathcal{A}$. We say that $\rho\in\mathcal{A}^{*}$ is positive if $\rho(x)\geq 0$ for any $x\in\mathcal{A}_{+}$, and $\mathcal{A}^{*}_{+}$ denotes the set of all positive linear functionals on $\mathcal{A}$. In addition, if $\rho(1_{\mathcal{A}})=1$ holds, then $\rho$ is said to be a state of $\mathcal{A}$. Here, $1_{\mathcal{A}}$ is the multiplicative unit in $\mathcal{A}$. We say that $\rho\in\mathcal{A}^{*}_{+}$ is faithful if $\rho(x)>0$ for any $x\in\mathcal{A}_{+}\backslash\{0\}$.
If $\mathcal{H}$ is finite-dimensional, then $B(\mathcal{H})$ coincides with $\mathrm{End}(\mathcal{H})$ (and $\mathbb{M}_{n}(\mathbb{C})$ if $\dim\mathcal{H}=n$). Namely, all linear operator on $\mathcal{H}$ are bounded. Moreover, the trace $\mathrm{Tr}$ on $B(\mathcal{H})$ can be defined. Using $\mathrm{Tr}$, we define the Hilbert–Schmidt inner product on $B(\mathcal{H})$ by $\langle A,B\rangle:=\mathrm{Tr}(B^{*}A)$ for all $A,B\in B(\mathcal{H})$. Moreover, by the Hilbert–Schmidt inner product, $B(\mathcal{H})$ can be naturally identified with its dual space $B(\mathcal{H})^{*}$.
We also use the following terminologies in convex analysis (see e.g., [11]). Let $K$ be a convex cone and $\leq$ denote the induced partial order on the vector space $E:=K-K$. A subset $\alpha\subset K$ is called a ray in $K$ if $\alpha=\{rx\mid r\geq 0\}$ for some $x\in K$. A ray $\alpha$ is said to be extreme if $ry\leq x$ implies $y\in\alpha$ for any $x\in\alpha$, $y\in K$, and $r\geq 0$. An element of extreme ray is said to be extreme in $K$, and $\mathrm{ex}(K)$ denotes the set of all extreme elements in $K$. By definition, $x\in\mathrm{ex}(K)$ holds if and only if $y\leq x$ implies $y=rx$ for some $r\geq 0$ for any $y\in K$. We suppose that $E$ is locally convex. We say that a nonnegative Borel measure $M$ on $K$ represents a point $x\in K$ such that $f(x)=\int_{K}f(y)M(dy)$ for any $f\in E^{*}$. For any two nonnegative Borel measures $M_{1},M_{2}$ on $K$ we write $M_{1}\prec M_{2}$ if $\int_{K}f(x)M_{1}(dx)\leq\int_{K}f(x)M_{2}(x)$ for any convex function $f$ on $K$.
We assume that $K$ is closed. A non-empty subset $C$ of $K$ is called a cap of $K$ if $C$ is compact, convex, and $K\backslash C$ is convex. See [11, p.80].
3. The generalized quantum de Finetti theorem via Choquet theory
Throughout this section, we fix two natural numbers $n,m$ and $\rho\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ that is faithful. The identity map on $\mathbb{M}_{n}(\mathbb{C})$ is simply denoted by $\mathrm{id}_{n}$. We define
$K_{m,\rho}:=\left\{(x_{l})_{l=0}^{\infty}\in\prod_{l=0}^{\infty}(\mathbb{M}_{m}(\mathbb{C})\otimes(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}})_{+}\,\middle|\,(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\rho)(x_{l+1})\leq x_{l}\text{ for all }l\geq 0\right\},$ (3.1)
where $\mathbb{M}_{n}(\mathbb{C})^{\otimes l}$ is the tensor product of $l$ copies of $\mathbb{M}_{n}(\mathbb{C})$ for $l\geq 1$, and $\mathbb{M}_{n}(\mathbb{C})^{\otimes 0}:=\mathbb{C}$. Moreover, $S_{l}$ is the symmetric group of degree $l$, and it acts on $\mathbb{M}_{n}(\mathbb{C})^{\otimes l}$ by
$\sigma\cdot(x_{1}\otimes\cdots\otimes x_{l}):=x_{\sigma(1)}\otimes\cdots\otimes x_{\sigma(l)}\quad(\sigma\in S_{l},\,x_{1},\dots,x_{l}\in\mathbb{M}_{n}(\mathbb{C})).$
We denote by $(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}}$ the *-subalgebra of fixed points under this action. We equip $K_{m,\rho}$ with the topology of component-wise weak convergence, where $(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}}$ is identified with the subspace of the $S_{l}$-invariant elements in $(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{*}$ by the Hilbert–Schmidt inner product. We also remark that $K_{m,\rho}$ is a convex cone by component-wise operations.
Lemma 3.1.
$K_{m,\rho}$ is the closed convex hull of $\mathrm{ex}(K_{m,\rho})$.
Proof.
For any $r>0$ we define
$C_{r}:=\{x=(x_{l})_{l=0}^{\infty}\in K_{m,\rho}\mid\mathrm{Tr}_{m}(x_{0})\leq r\}.$
For any $l\geq 1$ and $x\in C_{r}$ we have $(\mathrm{Tr}_{m}\otimes\rho^{\otimes l})(|x_{l}|)=(\mathrm{Tr}_{m}\otimes\rho^{\otimes l})(x_{l})\leq\mathrm{Tr}_{m}(x_{0})\leq 1$. Thus, $C_{r}$ is compact, and hence $C_{r}$ is a cap of $K_{m,\rho}$ (see [11, Proposition 13.2]). Moreover, $K_{m,\rho}=\bigcup_{r>0}C_{r}$ holds. Thus, the assertion follows from [11, Choquet theorem in p.81]. ∎
We remark that the above proof showed that $K_{m,\rho}$ is a union of its caps. Thus, combining with results in [11, Chapter 13], the following representation theorem holds true:
Proposition 3.1.
Every $x\in K_{m,\rho}\backslash\{0\}$ can be represented by a Borel probability measure $M_{x}$ on $K_{m,\rho}$ such that $M_{x}$ is maximal with respect to the partial order $\prec$ and supported on $\mathrm{ex}(K_{m,\rho})$ (i.e., $M_{x}(K_{m,\rho}\backslash\mathrm{ex}(K_{m,\rho}))=0$).
We obtain the following description of extreme points of $K_{m,\rho}$.
Theorem 3.1.
$\mathrm{ex}(K_{m,\rho})$ is contained in $\{(a\otimes b^{\otimes l})_{l=0}^{\infty}\mid a\in\mathbb{M}_{m}(\mathbb{C})_{+},b\in\mathbb{M}_{n}(\mathbb{C})_{+},\rho(b)\leq 1\}$.
Proof.
Let us fix $x=(x_{l})_{l=0}^{\infty}\in\mathrm{ex}(K_{m,\rho})$. For any $\nu\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ we define
$x_{\nu}=(x_{\nu,l})_{l=0}^{\infty}:=((\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\nu)(x_{l+1}))_{l=0}^{\infty}.$
Since $\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\nu$ is (completely) positive, we have $x_{\nu,l}\in(\mathbb{M}_{m}(\mathbb{C})\otimes(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}})_{+}$. Moreover, since $x_{\nu,l}$ is $S_{l}$-invariant, we have
$(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\rho)(x_{\nu,l+1})=(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\rho\otimes\nu)(x_{l+2})\leq(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\nu)(x_{l+1})=x_{\nu,l}.$
Thus, we have $x_{\nu}\in K_{m,\rho}$. If $\nu\leq\rho$, for any $l\geq 0$ we also have
$x_{\nu,l}=(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\nu)(x_{l+1})\leq(\mathrm{id}_{m}\otimes\mathrm{id}_{n}\otimes\rho)(x_{l+1})\leq x_{l}.$
Thus, $x_{\nu}\leq x$ holds, and hence there exists $b(\nu)\geq 0$ such that $x_{\nu}=b(\nu)x$ since $x\in\mathrm{ex}(K_{m,\rho})$. We claim that the mapping from $\nu$ to $b(\nu)$ can be extended linearly to a functional on $\mathbb{M}_{n}(\mathbb{C})^{*}$. Let $\nu\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ (not necessarily $\nu\leq\rho$). Since $\rho$ is faithful on $\mathbb{M}_{n}(\mathbb{C})$, there exists $r>0$ such that $r\nu\leq\rho$. Thus, for some $b(r\nu)>0$, we have $x_{r\nu}=b(r\nu)x$ and $rx_{\nu}=x_{r\nu}=b(r\nu)x$. Namely, $b(r\nu)/r$ depends only on $\nu$, and hence $b(\nu):=rb(\nu/r)$ is well defined. Moreover, the mapping $b\colon\mathbb{M}_{n}(\mathbb{C})^{*}_{+}\to\mathbb{R}_{\geq 0}$ is additive and positive-homogeneous. By the Jordan decomposition (see e.g., [10]), every linear functional on $\mathbb{M}_{n}(\mathbb{C})$ uniquely decomposes into a linear combination of four positive linear functionals. Thus, the mapping $b$ can be extended to a linear functional on $\mathbb{M}_{n}(\mathbb{C})$, and $b$ is naturally in $\mathbb{M}_{n}(\mathbb{C})=(\mathbb{M}_{n}(\mathbb{C})^{*})^{*}$. Moreover, we have $b\in\mathbb{M}_{n}(\mathbb{C})_{+}$, and $(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\nu)(x_{l+1})=(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\nu)(x_{l}\otimes b)$ for any $l\geq 0$ and $\nu\in\mathbb{M}_{n}(\mathbb{C})^{*}$. It implies that $x_{l+1}=x_{l}\otimes b$, and hence, iterating over $l$, we have $x_{l}=x_{0}\otimes b^{\otimes l}$. ∎
Let us recall that $a\in(\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C}))_{+}$ is called separable if $a$ is in the convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathbb{M}_{n}(\mathbb{C})_{+}$. Based on the results in this section, we obtain a characterization of separable elements, which is a generalization of the quantum de Finetti theorem.
Definition 3.1.
Let $\rho\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ and $a\in(\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C}))_{+}$. For every $l\geq 1$ we call $b$ in $(\mathbb{M}_{m}(\mathbb{C})\otimes(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}})_{+}$ a $l$-sub-extension of $a$ (with respect to $\rho$) if $(\mathrm{id}_{m}\otimes\mathrm{id}_{n}\otimes\rho^{\otimes(l-1)})(b)\leq a$ holds. If $a$ has a $l$-sub-extension, then $a$ is said to be $l$-sub-extendable (with respect to $\rho$).
Here we state our first main result.
Theorem 3.2 (generalized quantum de Finetti theorem).
For every $a\in(\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C}))_{+}$, the following are equivalent:
  1. (1)
    $a$ is separable.
  2. (2)
    There exists a faithful linear functional $\rho\in\mathbb{M}_{n}(\mathbb{C})_{+}^{*}$ such that $a$ is $l$-sub-extendable with respect to $\rho$ for every $l\geq 1$.
  3. (3)
    $a$ is $l$-sub-extendable with respect to every faithful linear functional on $\mathbb{M}_{n}(\mathbb{C})$ and $l\geq 1$.
Proof.
Let $\rho,\nu\in\mathbb{M}_{n}(\mathbb{C})_{+}^{*}$ be faithful. Then, there exists $r>0$ such that $r\nu\leq\rho$. Thus, if $b$ is a $l$-sub-extension of $a$ with respect to $\rho$, then $r^{l-1}b$ is a $l$-sub-extension of $a$ with respect to $\nu$. Namely, (2), (3) are equivalent.
We show that (1), (2) are equivalent. If $a$ is separable, then it is clearly $l$-sub-extendable for all $l\geq 1$. Conversely, we assume that there exist $l$-sub-extensions $x_{l}$ of $a$ for all $l\geq 1$. For any $k\geq l$ we define
$x_{l,k}:=(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\rho^{\otimes k-l})(x_{k})\in(\mathbb{M}_{m}(\mathbb{C})\otimes(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}})_{+},$
which is also $l$-sub-extension of $a$. Since $(\mathrm{Tr}_{m}\otimes\rho^{\otimes l})(|x_{l,k}|)=(\mathrm{Tr}_{m}\otimes\rho^{\otimes k})(x_{l,k})\leq(\mathrm{Tr}_{m}\otimes\rho)(a)$, the sequence $(x_{l,k})_{k\geq l}$ is contained in a compact subset of $(\mathbb{M}_{m}(\mathbb{C})\otimes(\mathbb{M}_{n}(\mathbb{C})^{\otimes l})^{S_{l}})_{+}$. Thus, there exists a convergence subsequence of $(x_{l,k})_{k\geq l}$, and $x_{l,\infty}$ denotes its limit. By applying this argument inductively, we obtain a sequence $(x_{l,\infty})_{l=2}^{\infty}$ of $l$-sub-extensions of $a$ such that $(\mathrm{id}_{m}\otimes\mathrm{id}_{n}^{\otimes l}\otimes\rho)(x_{l+1,\infty})\leq x_{l,\infty}$ for all $l\geq 2$. Namely, we have $(x_{l,\infty})_{l=0}^{\infty}\in K_{m,\rho}$, where $x_{1,\infty}=a$ and $x_{0,\infty}=(\mathrm{id}_{m}\otimes\rho)(a)$. By Lemma 3.1 and Theorem 3.1, $a$ can be approximated by convex combinations of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathbb{M}_{n}(\mathbb{C})_{+}$, and hence $a$ is separable. ∎
4. Matrix-valued Biane–Choquet–Deny theorem
4.1. *-Algebras from compact groups
We fix a compact group $G$. Let $C(G)$ denote the *-algebra of continuous functions on $G$. We denote by $\mathbb{C}[G]\subset C(G)$ the linear span of all matrix coefficients of finite-dimensional unitary representations. Here, by matrix coefficients, we mean continuous functions on $G$ given as $g\in G\mapsto\langle\pi(g)\xi,\eta\rangle$ for any (finite-dimensional) unitary representation $(\pi,\mathcal{H})$ of $G$ and two vectors $\xi,\eta\in\mathcal{H}$. By the definitions of tensor product and contragredient representations, $\mathbb{C}[G]$ becomes a *-subalgebra of $C(G)$. Moreover, $\mathbb{C}[G]$ is a Hopf *-algebra whose comultiplication $\Delta_{G}$, antipode $S_{G}$, and counit $\epsilon_{G}$ are given as follows:
  • the comultiplication $\Delta_{G}\colon\mathbb{C}[G]\to\mathbb{C}[G]\otimes\mathbb{C}[G]\subset C(G\times G)$ is the *-homomorphism given by $\Delta_{G}(f)(s,t):=f(st)$ for any $f\in\mathbb{C}[G]$ and $s,t\in G$
  • the antipode $S_{G}\colon\mathbb{C}[G]\to\mathbb{C}[G]$ is the anti-homomorphism given by $S_{G}(f)(s):=f(s^{-1})$ for any $f\in\mathbb{C}[G]$ and $s\in G$
  • the counit $\epsilon_{G}\colon\mathbb{C}[G]\to\mathbb{C}$ is the *-homomorphism given by $\epsilon_{G}(f):=f(e)$ for any $f\in\mathbb{C}[G]$, where $e\in G$ is the identity element.
[…]
$(xy,f):=(x\otimes y,\Delta_{G}(f)),\quad(x^{*},f):=\overline{(x,S_{G}(f)^{*})}\quad(x,y\in\mathcal{U}(G),\,f\in\mathbb{C}[G]).$
Then, $\epsilon_{G}$ becomes a multiplicative unit in $\mathcal{U}(G)$. In general, $\mathcal{U}(G)$ is not a Hopf *-algebra. However, a comultiplication-like unital *-homomorphism $\hat{\Delta}_{G}\colon\mathcal{U}(G)\to\mathcal{U}(G^{2})$ is defined by
$(\hat{\Delta}_{G}(x),f\otimes g):=(x,fg)\quad(x\in\mathcal{U}(G),\,f,g\in\mathbb{C}[G]).$
We remark that $\mathcal{U}(G)\otimes\mathcal{U}(G)$ can be naturally embedded into $\mathcal{U}(G^{2})$, but $\mathcal{U}(G^{2})$ is strictly larger than $\mathcal{U}(G)\otimes\mathcal{U}(G)$ in general. Similarly, two unital *-homomorphisms $\hat{\Delta}_{G}\otimes\mathrm{id}$ and $\mathrm{id}\otimes\hat{\Delta}_{G}$ from $\mathcal{U}(G^{2})$ to $\mathcal{U}(G^{3})$ are defined by
$((\hat{\Delta}_{G}\otimes\mathrm{id})(x),f\otimes g\otimes h):=(x,fg\otimes h),\quad((\mathrm{id}\otimes\hat{\Delta}_{G})(x),f\otimes g\otimes h):=(x,f\otimes gh)$
for any $x\in\mathcal{U}(G^{2})$ and $f,g,h\in\mathbb{C}[G]$. By the associativity of the multiplication of $\mathbb{C}[G]$, we have $(\hat{\Delta}_{G}\otimes\mathrm{id})\circ\hat{\Delta}_{G}=(\mathrm{id}\otimes\hat{\Delta}_{G})\circ\hat{\Delta}_{G}$, and this is denoted by $\hat{\Delta}_{G}^{(2)}$. Inductively, we also obtain a unital *-homomorphism $\hat{\Delta}_{G}^{(l)}\colon\mathcal{U}(G)\to\mathcal{U}(G^{l})$ for any $l\geq 2$.
Let $(\pi,\mathcal{H})$ be a finite-dimensional unitary representation of $G$. For any $\xi,\eta\in\mathcal{H}$ we denote by $c^{\pi}_{\xi,\eta}$ the corresponding matrix coefficient, that is, $c^{\pi}_{\xi,\eta}(t):=\langle\pi(t)\xi,\eta\rangle$. For any $x\in\mathcal{U}(G)$ we define $\pi(x)\in B(\mathcal{H})$ as the unique linear operator satisfying $\langle\pi(x)\xi,\eta\rangle=(x,c^{\pi}_{\xi,\eta})$. Then, it is easy to show that the mapping $x\in\mathcal{U}(G)\mapsto\pi(x)\in B(\mathcal{H})$ is a unital *-homomorphism. We remark that $\pi(e_{t})=\pi(t)$ for any $t\in G$, where $e_{t}\in\mathcal{U}(G)$ is the evaluation map at $t$, that is, $(e_{t},f):=f(t)$ for any $f\in\mathbb{C}[G]$. By the definition of the *-operation on $\mathcal{U}(G)$, we have $e_{t}^{*}=e_{t^{-1}}$ for any $t\in G$. Thus, any *-representation of $\mathcal{U}(G)$ produces a unitary representation of $G$. Similarly, for any two unitary representations $(\pi,\mathcal{H})$ and $(\lambda,\mathcal{K})$ of $G$ we obtain a unital *-homomorphism $\pi\otimes\lambda\colon\mathcal{U}(G^{2})\to B(\mathcal{H}\otimes\mathcal{K})$ by $\langle(\pi\otimes\lambda)(x)\xi_{1}\otimes\xi_{2},\eta_{1}\otimes\eta_{2}\rangle=(x,c^{\pi}_{\xi_{1},\eta_{1}}\otimes c^{\lambda}_{\xi_{2},\eta_{2}})$ for any vectors $\xi_{1},\eta_{1}\in\mathcal{H}$ and $\xi_{2},\eta_{2}\in\mathcal{K}$. Namely, the representation of $\mathcal{U}(G)$ given as $(\pi\otimes\lambda)\circ\hat{\Delta}_{G}$ corresponds to the tensor product representation $(\pi\otimes\lambda,\mathcal{H}\otimes\mathcal{K})$ of $G$. In what follows, $(\pi\otimes\lambda)\circ\hat{\Delta}_{G}$ is simply denoted by $\pi\otimes\lambda$.
Let $\widehat{G}$ denote the set of all equivalence classes of irreducible representations of $G$, and it is always assumed to be countable. Throughout the paper, we fix representatives $((\pi_{\alpha},\mathcal{H}_{\alpha}))_{\alpha\in\widehat{G}}$ of all irreducible representations. By the Peter–Weyl theorem, $\mathcal{U}(G)$ is isomorphic to $\prod_{\alpha\in\widehat{G}}B(\mathcal{H}_{\alpha})$ as a *-algebra, and the isomorphism is given by $\Pi_{G}(x):=(\pi_{\alpha}(x))_{\alpha\in\widehat{G}}$. In what follows, we freely identify $\mathcal{U}(G)$ and $\prod_{\alpha\in\widehat{G}}B(\mathcal{H}_{\alpha})$.
4.2. Matrix-valued Martin boundaries on the duals of compact groups
First, we mention the perspective of quantum random walks or random walks on non-commutative spaces. See [3] and references therein for more details. From this standpoint, a pair consisting of a unital *-algebra and a state is interpreted as a non-commutative analog of a probability space. Specially, a unital *-algebra serves as an algebra of random variables, while a state represents a linear functional obtained through expectation. Thus, a state is a non-commutative counterpart of a probability measure. Our results in this section concern the non-commutative probability space from $\mathcal{U}(G)$. See also the comment below (4.1).
We remark that $x\in\mathcal{U}(G)_{+}$ if and only if $\pi(x)\in B(\mathcal{H})_{+}$ for any finite-dimensional unitary representation $(\pi,\mathcal{H})$ of $G$. We also say that $\rho\in\mathcal{U}(G)^{*}$ is positive if $\rho(x)\geq 0$ for any $x\in\mathcal{U}(G)_{+}$. By the dual pairing between $\mathcal{U}(G)$ and $\mathbb{C}[G]$, every $f\in\mathbb{C}[G]$ is naturally identified with $\widehat{f}\in\mathcal{U}(G)^{*}$ given by $\widehat{f}(x):=(x,f)$ for any $x\in\mathcal{U}(G)$. We denote by $\mathcal{U}(G)_{*}$ the image of this identification, i.e., $\mathcal{U}(G)_{*}:=\{\widehat{f}\mid f\in\mathbb{C}[G]\}$ and $\mathcal{U}(G)_{*,+}:=\mathcal{U}(G)_{*}\cap\mathcal{U}(G)^{*}_{+}$. We will use the following notation:
Definition 4.1.
For any $\rho\in\mathcal{U}(G)_{*}$ we define $f_{\rho}\in\mathbb{C}[G]$ by $\widehat{f_{\rho}}=\rho$.
For any finite-dimensional unitary representation $(\pi,\mathcal{K})$ of $G$ and any $\xi\in\mathcal{K}$, let $\omega_{\xi}$ denote the positive linear form on $B(\mathcal{K})$ given by $\omega_{\xi}(x):=\langle x\xi,\xi\rangle$ for any $x\in B(\mathcal{K})$. Then $\omega_{\xi}\circ\pi\in\mathcal{U}(G)_{*,+}$ and moreover any element $\rho\in\mathcal{U}(G)_{*,+}$ has this form. Let $\rho\in\mathcal{U}(G)_{*,+}$ be given. Note first that by Schur orthogonality relations, $\rho$ is finitely supported, i.e., $\rho|_{B(\mathcal{H}_{\alpha})}=0$ holds for any but finitely many $\alpha\in\widehat{G}$. For $\rho|_{B(\mathcal{H}_{\alpha})}$, let $(\sigma_{\alpha},\mathcal{K}_{\alpha})$ be the GNS representation of $B(\mathcal{H}_{\alpha})$ and let $\xi_{\alpha}\in\mathcal{K}_{\alpha}$ such that $\rho|_{B(\mathcal{H}_{\alpha})}(x)=\langle\sigma_{\alpha}(x)\xi_{\alpha},\xi_{\alpha}\rangle$ for any $x\in B(\mathcal{H}_{\alpha})$. Note that $\mathcal{K}_{\alpha}$ is finite-dimensional, and $\mathcal{K}_{\alpha}=\{0\}$ for all but finitely many $\alpha\in\widehat{G}$. Let $\sigma$ be the direct sum
$\sigma:=\bigoplus_{\alpha\in\widehat{G}}\sigma_{\alpha}\colon\mathcal{U}(G)\to\bigoplus_{\alpha\in\widehat{G}}B(\mathcal{K}_{\alpha}).$
It is a finite-dimensional representation, and we have $\rho(x)=\langle\sigma(x)\xi,\xi\rangle$ for any $x\in\mathcal{U}(G)$, where $\xi:=\sum_{\alpha\in\widehat{G}}\xi_{\alpha}$.
For any $\rho\in\mathcal{U}(G)_{*}$ we define a linear map $P_{\rho}\colon\mathcal{U}(G)\to\mathcal{U}(G)$ by “$P_{\rho}:=(\mathrm{id}\otimes\rho)\circ\Delta_{G}$”. The precise definition is as follows:
Definition 4.2.
For any $\rho\in\mathcal{U}(G)_{*}$ we define a linear map $P_{\rho}\colon\mathcal{U}(G)\to\mathcal{U}(G)$ by
$(P_{\rho}(x),f):=(\hat{\Delta}_{G}(x),f\otimes f_{\rho})=(x,ff_{\rho})\quad(x\in\mathcal{U}(G),\,f\in\mathbb{C}[G]).$
Its matrix-valued version is also defined as $P_{m,\rho}:=\mathrm{id}_{m}\otimes\rho$ on $\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G)$.
For any $\rho,\nu\in\mathcal{U}(G)_{*}$ the convolution $\rho*\nu\in\mathcal{U}(G)_{*}$ is defined as $\rho*\nu:=\rho\circ P_{\nu}$. It is easy to check $\rho*\nu=(\rho\otimes\nu)\circ\hat{\Delta}_{G}$, where $\rho\otimes\nu$ is a linear functional naturally defined on $\mathcal{U}(G^{2})$ for $\rho,\nu\in\mathcal{U}(G)_{*}$. Since $\mathbb{C}[G]$ is commutative, we have $\rho*\nu=\nu*\rho$, and equivalently, $P_{\rho}\circ P_{\nu}=P_{\nu}\circ P_{\rho}$ holds. We also define $\rho^{*l}:=\rho\circ P_{\rho}^{l-1}$ for any $l\geq 1$ and $\rho^{*0}:=\epsilon_{G}$.
Proposition 4.1.
For any $\rho\in\mathcal{U}(G)_{*,+}$ the linear map $P_{\rho}$ is completely positive, that is, $P_{m,\rho}$ is positivity-preserving for all $m\geq 0$.
Proof.
We may assume that $\rho=\omega_{\xi}\circ\pi$ for some finite-dimensional unitary representation $(\pi,\mathcal{K})$ of $G$ and $\xi\in\mathcal{K}$. Let $x\in\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G)$ be positive. Our goal is to show $P_{m,\rho}(x)\geq 0$. For any finite-dimensional unitary representation $(\lambda,\mathcal{H})$ of $G$ we have
$(\mathrm{id}_{m}\otimes\lambda)(P_{m,\rho}(x))=(\mathrm{id}_{m}\otimes\mathrm{id}_{B(\mathcal{H})}\otimes\omega_{\xi})(\mathrm{id}_{m}\otimes\lambda\otimes\pi)(x)\geq 0,$
that is, $P_{m,\rho}(x)\geq 0$. ∎
We define
$B_{m,\rho}:=\{x\in(\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G))_{+}\mid P_{m,\rho}(x)\leq x\}.$ (4.1)
In the viewpoint of noncommutative probability theory, $P_{\rho}$ is regarded as the transition operator of a Markov process. Thus, $B_{m,\rho}$ can be naturally interpreted as the space of $\mathbb{M}_{m}(\mathbb{C})$-valued $P_{\rho}$-subharmonic functions. In the paper, an element in $B_{m,\rho}$ is called $P_{m,\rho}$-subharmonic. If $P_{m,\rho}(x)=x$, then $x$ is called $P_{m,\rho}$-harmonic.
We equip $\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G)$ with the weak-topology induced by $\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{C}[G]$. Here, $\mathbb{M}_{m}(\mathbb{C})$ is identified with its dual space by the Hilbert–Schmidt inner product. By $\mathrm{id}_{m}\otimes\Pi_{G}$, we have $\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G)\cong\prod_{\alpha\in\widehat{G}}\mathbb{M}_{m}(\mathbb{C})\otimes B(\mathcal{H}_{\alpha})$, where the right-hand side is endowed with the product topology of the weak topologies, and this is equivalent to the topology of $\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G)$.
By a proof as in Lemma 3.1, we have
Lemma 4.1.
$B_{m,\rho}$ is the closed convex hull of $\mathrm{ex}(B_{m,\rho})$.
Proof.
For any $c=(c_{\alpha})_{\alpha\in\widehat{G}}\in\mathbb{R}_{>0}^{\widehat{G}}$ we define $\varphi_{c}\colon(\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G))_{+}\to[0,\infty]$ by
$\varphi_{c}(x)=\sum_{\alpha\in\widehat{G}}c_{\alpha}(\mathrm{Tr}_{m}\otimes\mathrm{Tr}_{\alpha}\circ\pi_{\alpha})(x),$
where $\mathrm{Tr}_{\alpha}$ is the non-normalized trace on $B(\mathcal{H}_{\alpha})$. Then, $\varphi_{c}$ is lower semi-continuos, additive, and positive-homogeneous. Moreover, $C_{\varphi_{c}}:=\{x\in B_{m,\rho}\mid\varphi_{c}(x)\leq 1\}$ is compact since
$C_{\varphi_{c}}\subset\{x\in\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G)\mid(\mathrm{Tr}_{m}\otimes\mathrm{Tr}_{\alpha}\circ\pi_{\alpha})(|x|)<c_{\alpha}^{-1}\text{ for any }\alpha\in\widehat{G}\}.$
Thus, $C_{\varphi_{c}}$ is a cap of $B_{m,\rho}$ (see [11, Proposition 13.2]). Thus, the claim follows from [11, Chqoeut theorem in p.81] since $B_{m,\rho}=\bigcup_{c\in\mathbb{R}_{>0}^{\widehat{G}}}C_{\varphi_{c}}$. ∎
As in Lemma 3.1, we showed that $B_{m,\rho}$ is a union of its caps in the above proof. Thus, combining with results in [11, Chapter 13], we obtain the following representation theorem:
Proposition 4.2.
Every $x\in B_{m,\rho}\backslash\{0\}$ can be represented by a maximal Borel probability measure $M_{x}$ supported on $\mathrm{ex}(B_{m,\rho})$ (in the sense of Proposition 3.1).
We now study the extreme points $\mathrm{ex}(B_{m,\rho})$ of $B_{m,\rho}$ and obtain a matrix-valued extension of a theorem by Biane [1] that is itself a non-commutative version of the Choquet–Deny theorem.
[…]
$(P_{\rho}(e),f)=(\hat{\Delta}_{G}(e),f\otimes f_{\rho})=\rho(e)(e,f).$
Thus, an exponential element $e\in\mathcal{E}(G)$ is $P_{\rho}$-subharmonic if and only if $\rho(e)\leq 1$. If $\mathcal{E}_{\rho}(G)$ denotes the intersection $\mathcal{E}(G)\cap B_{1,\rho}$, we then have
$\mathcal{E}_{\rho}(G)=\{e\in\mathcal{E}(G)\mid\rho(e)\leq 1\}.$
We assume that $\rho\in\mathcal{U}(G)_{*,+}$ is generating, i.e., for every $\nu\in\mathcal{U}(G)_{*,+}$ we have $\nu\leq\sum_{l=0}^{k}a_{l}\rho^{*l}$ for some $a_{l}\geq 0$ ($l=0,\dots,k$) and $k\geq 0$.
Theorem 4.1 (matrix-valued Biane–Choquet–Deny theorem).
The set $\mathrm{ex}(B_{m,\rho})$ of extreme points of $B_{m,\rho}$ is contained in $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$. In particular, $B_{m,\rho}$ is the closed convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$.
Proof.
Let $x\in\mathrm{ex}(B_{m,\rho})$. For any $\nu\in\mathcal{U}(G)_{*,+}$ we have $P_{m,\nu}(x)\geq 0$ by Proposition 4.1. Moreover, $P_{m,\rho}(P_{m,\nu}(x))=P_{m,\nu}(P_{m,\rho}(x))\leq P_{m,\nu}(x)$, and hence $P_{m,\nu}(x)\in B_{m,\rho}$. Since $\rho$ is generating, we have $\nu\leq\sum_{l=0}^{k}a_{l}\rho^{*l}$ for some $a_{l}\geq 0$ ($l=0,\dots,k$) and $k\geq 0$. Thus, we have
$P_{m,\nu}(x)\leq\sum_{l=0}^{k}a_{l}P_{\rho}^{l}(x)\leq\left(\sum_{l=0}^{k}a_{l}\right)x,$
and hence there exists $\alpha(\nu)\geq 0$ such that $P_{m,\nu}(x)=\alpha(\nu)x$. By the Jordan decomposition (see e.g., [10]), any $\nu\in\mathcal{U}(G)_{*}$ is uniquely decomposed into a summation of four elements in $\mathcal{U}(G)_{*,+}$. Thus, $\alpha$ induces a linear functional on $\mathcal{U}(G)_{*}\cong\mathbb{C}[G]$. Therefore, there exists $e\in\mathcal{U}(G)_{+}$ such that $\nu(e)=\alpha(\nu)$ for any $\nu\in\mathcal{U}(G)$. Namely, we have $P_{m,\nu}(x)=(e,f_{\nu})x$.
For any $\mu\in\mathbb{M}_{m}(\mathbb{C})^{*}$, we define $x_{\mu}:=(\mu\otimes\mathrm{id}_{\mathcal{U}(G)})(x)\in\mathcal{U}(G)$. For any $f,g\in\mathbb{C}[G]$ we have
$(\hat{\Delta}_{G}(x_{\mu}),f\otimes g)=(P_{\widehat{f}}(x_{\mu}),g)=(e\otimes x_{\mu},f\otimes g).$
Moreover, since $fg=gf$, we have
$(\hat{\Delta}_{G}(x_{\mu}),f\otimes g)=(P_{\widehat{g}}(x_{\mu}),f)=(x_{\mu}\otimes e,f\otimes g).$
Thus, $\hat{\Delta}_{G}(x_{\mu})=x_{\mu}\otimes e=e\otimes x_{\mu}$ holds, and $e$ is group-like. Moreover, there exists $a\in\mathbb{M}_{m}(\mathbb{C})_{+}$ such that $x=a\otimes e$. Since $x$ is $P_{m,\rho}$-subharmonic, $e$ is also $P_{\rho}$-subharmonic, i.e., $e\in\mathcal{E}_{\rho}(G)$. Therefore, we have $\mathrm{ex}(B_{m,\rho})\subset\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$, and hence $B_{m,\rho}$ is contained in the closed convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$. On the other hand, since $\mathcal{E}_{\rho}(G)\subset B_{1,\rho}$ by definition, we have $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)\subset B_{m,\rho}$. Thus, the closed convex hull of $\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(G)$ is contained in $B_{m,\rho}$, that is, they are equal. ∎
Remark 4.1.
Our results in this section can be extended to compact quantum groups with slight modifications. See, e.g., [9] for the basics of compact quantum groups. If $G$ is a compact quantum group, then $\mathbb{C}[G]$ is non-commutative, and hence the convolution product on $\mathcal{U}(G)_{*}$ is also non-commutative. However, for any $\rho\in\mathcal{U}(G)_{*,+}$, Propositions 4.1, 4.2 and Lemma 4.1 hold true. In addition, if $\rho$ is generating and central (i.e., $\rho$ satisfies $\rho*\nu=\nu*\rho$ for any $\nu\in\mathcal{U}(G)_{*}$), then Theorem 4.1 holds true. However, we note that, in the case of $SU_{q}(2)$ for instance, every central element in $\mathcal{U}(SU_{q}(2))_{*}$ is proportional to the counit (i.e., it is not generating) since the center of $\mathbb{C}[SU_{q}(2)]$ is trivial (see [8, Corollary 4.4]).
By Theorem 4.1, we can realize that $B_{m,\rho}$ is similar to the convex cones of separable elements. Namely, any subharmonic elements are produced from tensors of two positive elements. This observation and the generalized quantum de Finetti theorem (see Theorem 3.2) suggest the relation between $B_{m,\rho}$ and sub-extendability. We will investigate this point in the next section.
5. The relation between the matrix-valued Martin boundary and the generalized quantum de Finetti theorem: quantum random walks
In this part, we deepen our understanding of the relation between matrix-valued Martin boundary and the generalized quantum de Finetti theorem. Throughout this section, we fix a finite-dimensional unitary representation $(\Pi,\mathcal{H})$ of $G$ and $\rho\in B(\mathcal{H})_{*,+}$ that is faithful. Let us denote by the same symbol $\rho$ the linear functional on $\mathcal{U}(G)$ defined as $\rho\circ\Pi$. We remark that $\rho$ is generating if $(\Pi,\mathcal{H})$ is also generating $\widehat{G}$, that is, for any $\alpha\in\widehat{G}$ there exists $l\geq 0$ such that $(\pi_{\alpha},\mathcal{H}_{\alpha})$ is a sub-representation of $(\Pi^{\otimes l},\mathcal{H}^{\otimes l})$.
Let us recall that the UHF (uniformly hyperfinite) algebra $B(\mathcal{H})^{\otimes\infty}$ is defined as the $C^{*}$-inductive limit of the $B(\mathcal{H})^{\otimes l}$ ($l\geq 0$). If $\rho$ is a state of $B(\mathcal{H})$, then the infinite tensor product state $\rho^{\otimes\infty}$ on $B(\mathcal{H})^{\otimes\infty}$ is defined. Thus, we obtain the $C^{*}$-probability space $(B(\mathcal{H})^{\otimes\infty},\rho^{\otimes\infty})$. Moreover, we can regard the family of the subalgebras $B(\mathcal{H})^{\otimes l}$ as a filtration of $\sigma$-algebras. Namely, $(B(\mathcal{H})^{\otimes\infty},(B(\mathcal{H})^{\otimes l})_{l=0}^{\infty},\rho^{\otimes\infty})$ is a filtered $C^{*}$-probability space. For every $l\geq 0$ the conditional expectation $E_{l}$ from $B(\mathcal{H})^{\otimes\infty}$ to $B(\mathcal{H})^{\otimes l}$ is also defined as the linear map induced by $\mathrm{id}_{B(\mathcal{H})}^{\otimes l}\otimes\rho^{\otimes k}\colon B(\mathcal{H})^{\otimes l+k}\to B(\mathcal{H})^{\otimes l}$ ($k\geq 0$). Under this point of view, a sequence $(x_{l})_{l=0}^{\infty}$ in $(B(\mathcal{H})^{\otimes\infty},\rho^{\otimes\infty})$ is said to be adapted sub-martingale if $x_{l}\in B(\mathcal{H})^{\otimes l}$ and $E_{l}(x_{l+1})\leq x_{l}$ for every $l\geq 0$. According to (3.1), we can interpret $K_{m,\rho}$ as the set of paths of $\mathbb{M}_{m}(\mathbb{C})$-valued adapted sub-martingale processes. Here, $B(\mathcal{H})$ is naturally identified with $\mathbb{M}_{n}(\mathbb{C})$ if $\dim\mathcal{H}=n$. Such an interpretation was given initially by Izumi [6]. In his work, he initiated the non-commutative Poisson boundary theory. In the paper, our setting is slightly different from his work, but we can find a connection between such a non-commutative boundary theory and quantum information theory.
We compare $K_{m,\rho}$ and $B_{m,\rho}$. For any $l\geq 0$ we define
$\Pi_{m,l}:=(\mathrm{id}_{m}\otimes\Pi^{\otimes l})\colon\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G)\to\mathbb{M}_{m}(\mathbb{C})\otimes B(\mathcal{H})^{\otimes l},$
$\Pi_{m,\infty}\colon\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G)\to\prod_{l=0}^{\infty}\mathbb{M}_{m}(\mathbb{C})\otimes B(\mathcal{H})^{\otimes l}$
by $\Pi_{m,\infty}(x)=(\Pi_{m,l}(x))_{l=0}^{\infty}$, where $\Pi^{\otimes 0}:=\epsilon_{G}$. We remark that $\Pi_{m,l}(x)\in\mathbb{M}_{m}(\mathbb{C})\otimes(B(\mathcal{H})^{\otimes l})^{S_{l}}$ for any $x\in\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G)$ and $l\geq 0$. In fact, we may observe the following: Let $\mu\in\mathbb{M}_{m}(\mathbb{C})^{*}$ and $x_{\mu}:=(\mu\otimes\mathrm{id}_{\mathcal{U}(G)})(x)\in\mathcal{U}(G)$. Since $\mathbb{C}[G]$ is a commutative algebra, for any $\sigma\in S_{l}$ and $\nu_{1},\dots,\nu_{l}\in B(\mathcal{H})^{*}$ and $\mu\in\mathbb{M}_{m}(\mathbb{C})^{*}$ we have
$(\nu_{1}\otimes\cdots\otimes\nu_{l})(\sigma\cdot\Pi_{1,l}(x_{\mu}))$ $=(\nu_{\sigma^{-1}(1)}\otimes\cdots\otimes\nu_{\sigma^{-1}(l)})(\Pi_{1,l}(x_{\mu}))$
$=(\hat{\Delta}_{G}^{(l)}(x_{\mu}),f_{\nu_{\sigma^{-1}(1)}\circ\Pi}\otimes\cdots\otimes f_{\nu_{\sigma^{-1}(l)}\circ\Pi})$
$=(x_{\mu},f_{\nu_{1}\circ\Pi}\cdots f_{\nu_{l}\circ\Pi})$
$=(\nu_{1}\otimes\cdots\otimes\nu_{l})(\Pi_{1,l}(x_{\mu})).$
Thus, we have $\sigma\cdot\Pi_{1,l}(x_{\mu})=\Pi_{1,l}(x_{\mu})$, and hence $\Pi_{m,l}(x)\in\mathbb{M}_{m}(\mathbb{C})\otimes(B(\mathcal{H})^{\otimes l})^{S_{l}}$.
We start to discuss the relation among the results of the previous two sections.
Proposition 5.1.
$\Pi_{m,\infty}(B_{m,\rho})$ is contained in $K_{m,\rho}$. Moreover, $\Pi_{m,\infty}$ is injective if $(\Pi,\mathcal{H})$ is generating $\widehat{G}$.
Proof.
Ler $x\in B_{m,\rho}$. For any $l\geq 0$ we already know that $\Pi_{m,l}(x)$ is in $(\mathbb{M}_{m}(\mathbb{C})\otimes(B(\mathcal{H})^{\otimes l})^{S_{l}})_{+}$. Moreover, for any $\nu\in(B(\mathcal{H})^{\otimes l})^{*}_{+}$ we have
$(\mathrm{id}_{m}\otimes\nu\otimes\rho)(\Pi_{m,l+1}(x))$ $=(\mathrm{id}_{m}\otimes((\nu\circ\Pi^{\otimes l})*\rho))(x)$
$=(\mathrm{id}_{m}\otimes\nu\circ\Pi^{\otimes l})(P_{m,\rho}(x))$
$\leq(\mathrm{id}_{m}\otimes\nu\circ\Pi^{\otimes l})(x)$
$=(\mathrm{id}_{m}\otimes\nu)(\Pi_{m,l}(x)).$
Therefore, $(\mathrm{id}_{m}\otimes\mathrm{id}_{B(\mathcal{H})}^{\otimes l}\otimes\rho)(\Pi_{m,l+1}(x))\leq\Pi_{m,l}(x)$ holds, and hence $\Pi_{m,\infty}(x)\in K_{m,\rho}$.
We assume that $(\Pi,\mathcal{H})$ is generating $\widehat{G}$ and prove that $\Pi_{m,\infty}$ is injective. We suppose that $\Pi_{m,\infty}(x)=\Pi_{m,\infty}(y)$ for $x,y\in\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G)$. Since $(\Pi,\mathcal{H})$ is generating $\widehat{G}$, for any $\alpha\in\widehat{G}$ the corresponding irreducible representation $(\pi_{\alpha},\mathcal{H}_{\alpha})$ is contained in $(\Pi^{\otimes l},\mathcal{H}^{\otimes l})$ for some $l\geq 0$. Since $\Pi_{m,l}(x)=\Pi_{m,l}(y)$, we have $(\mathrm{id}_{m}\otimes\pi_{\alpha})(x)=(\mathrm{id}_{m}\otimes\pi_{\alpha})(y)$. Thus, $x=y$ holds, and hence $\Pi_{m,\infty}$ is injective. ∎
The next statement immediately follows from Theorem 4.1 if $(\Pi,\mathcal{H})$ is generating $\widehat{G}$. However, by Lemma 3.1, Theorem 3.1, and the above proposition, the following holds true even if $(\Pi,\mathcal{H})$ is not generating $\widehat{G}$.
Corollary 5.1.
$\Pi_{m,1}(B_{m,\rho})$ is contained in the set of separable elements in $\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C})$.
6. The case $G=U(n)$ and the generalized quantum de Finetti theorem revisited
In this section, we consider the case $G=U(n)$. For any $t\in U(n)$, recall that $e_{t}\in\mathcal{U}(U(n))$ denotes the evaluation at $t$, i.e., $(e_{t},f)=f(t)$ for all $f\in\mathbb{C}[U(n)]$. It is a group-like of $\mathcal{U}(U(n))$, as for any $f,g\in\mathbb{C}[U(n)]$
$(\hat{\Delta}_{U(n)}(e_{t}),f\otimes g)=(e_{t},fg)=f(t)g(t)=(e_{t}\otimes e_{t},f\otimes g),$
By Weyl’s unitary trick, every finite-dimensional representation of $U(n)$ can be extended to a representation of $GL(n,\mathbb{C})$. Thus, for every $t\in GL(n,\mathbb{C})$ its evaluation map $e_{t}$ on $\mathbb{C}[U(n)]$ is well defined, and is also group-like.
The following was initially mentioned by Biane [2, Theorem 5.1] when $G$ is a connected, simply connected semisimple compact Lie group. Using the same argument, we can then show for $G=U(n)$.
Lemma 6.1.
The mapping $t\in GL(n,\mathbb{C})\mapsto e_{t}\in\mathcal{G}(U(n))$ is a bijection. Moreover, $e_{t}\in\mathcal{E}(U(n))$ if and only if $t\in GL(n,\mathbb{C})_{+}:=GL(n,\mathbb{C})\cap\mathbb{M}_{n}(\mathbb{C})_{+}$.
Proof.
Let $e\in\mathcal{G}(U(n))$. By Weyl’s unitary trick, $e$ gives a homomorphism from $\mathbb{C}[GL(n,\mathbb{C})]$ to $\mathbb{C}$. We remark that $GL(n,\mathbb{C})$ is an algebraic set by the following identification:
$GL(n,\mathbb{C})\cong\{((x_{ij})_{i,j=1}^{n},y)\in\mathbb{C}^{n^{2}+1}\mid y\det(x_{ij})_{i,j=1}^{n}-1=0\},$
where the identification map is given as $x\mapsto(x,\det(x)^{-1})$. Thus, by Hilbert’s nullstellensatz, every maximal ideal of $\mathbb{C}[GL(n,\mathbb{C})]$ is the kernel of an evaluation map. In particular, $\ker(e)$ is a maximal ideal, and hence $\ker(e)=\ker(e_{t})$ for some $t\in GL(n,\mathbb{C})$. Therefore, we have $e=e_{t}$, and the mapping $t\mapsto e_{t}$ is a bijection from $GL(n,\mathbb{C})$ to $\mathcal{G}(U(n))$.
If $e_{t}$ is positive, then so is the fundamental representation of $e_{t}$, i.e., $t$ itself is positive. Conversely, if $t\in GL(n,\mathbb{C})_{+}$, then we consider its diagonalization $t=u\mathrm{diag}(a_{1},\dots,a_{n})u^{*}$, where $u\in U(n)$ and $\mathrm{diag}(a_{1},\dots,a_{n})$ is the diagonal matrix with entries $a_{1},\dots,a_{n}\geq 0$. For any unitary representation $(\pi,\mathcal{H})$ of $U(n)$,
$\pi(e_{t})=\pi(t)=\pi(u)\pi(\mathrm{diag}(a_{1},\dots,a_{n}))\pi(u)^{*}$
holds, where we remark that $\pi(u)$ is unitary. Moreover, it is well known that $\mathcal{H}$ is decomposed into a direct sum of root spaces. For any $(\lambda_{1},\dots,\lambda_{n})\in\mathbb{Z}^{n}$ the diagonal part $\pi(\mathrm{diag}(a_{1},\dots,a_{n}))$ acts as the scalar $a_{1}^{\lambda_{1}}\cdots a_{n}^{\lambda_{n}}$ on the corresponding root space. Thus, $\pi(\mathrm{diag}(a_{1},\dots,a_{n}))$ is positive, and hence so is $\pi(e_{t})$. Namely, $e_{t}$ is exponential. ∎
Remark 6.1.
The description of group-like elements (i.e., the first statement of Lemma 6.1) is given in [9, Theorem 3.2.2] for arbitrary compact Lie group.
Remark 6.2.
An alternative proof of Lemma 6.1 is also given as follows. Let us consider the embedding $U(n)\hookrightarrow SU(n+1)$ by $u\mapsto u\oplus\det(u)^{-1}$. Thus, we obtain the restriction map from $\mathbb{C}[SU(n+1)]$ to $\mathbb{C}[U(n)]$. Then, its dual map $\mathcal{U}(U(n))\to\mathcal{U}(SU(n+1))$ is an injective unital *-homomorphism intertwining both comultiplications. Thus, if $e$ is a group-like element in $\mathcal{U}(U(n))$, then so is in $\mathcal{U}(SU(n+1))$. By [2, Theorem 5.1], there exists $t\in SL(n+1,\mathbb{C})$ such that $e=e_{t}$. Let us consider the representation of $\mathcal{U}(SU(n+1))$ induced by the fundamental representation of $SU(n+1)$. Since $e\in\mathcal{U}(U(n))$, the image of $e$ (i.e., $t$) has the form $t^{\prime}\oplus\det(t^{\prime})^{-1}$ for some $t^{\prime}\in GL(n,\mathbb{C})$. Therefore, $e=e_{t^{\prime}}$ as elements in $\mathcal{U}(U(n))$. Moreover, by [2, Theorem 5.1], $e$ is positive if and only if $t\in SL(n+1,\mathbb{C})_{+}$, and it is equivalent to that $t^{\prime}\in GL(n,\mathbb{C})_{+}$.
In what follows, let $(\Pi,\mathbb{C}^{n})$ denote the fundamental representation of $U(n)$ and fix a faithful positive linear functional $\rho$ on $\mathbb{M}_{n}(\mathbb{C})$ ($=B(\mathbb{C}^{n})$). As in the previous section, $\rho$ will also denote the linear functional $\rho\circ\Pi$ on $\mathcal{U}(U(n))$. We also define $\Pi_{m,l}$ and $\Pi_{m,\infty}$ as the same as in the previous section. We remark that even if the fundamental representation is not generating $\widehat{U(n)}$, the following statement holds true (see also Proposition 5.1).
Proposition 6.1.
$\Pi_{m,\infty}$ is injective on $\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{G}(U(n))$.
Proof.
We suppose that $\Pi_{m,\infty}(e_{1})=\Pi_{m,\infty}(e_{2})$ for $e_{1},e_{2}\in\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{G}(U(n))$. In the representation theory, it is well known that $\widehat{U(n)}$ can be parametrized by $\{\lambda=(\lambda_{1}\geq\cdots\geq\lambda_{n})\in\mathbb{Z}^{n}\}$. By the Schur–Weyl duality, for any $\lambda$ with $\lambda_{n}\geq 0$ the associated irreducible representation, denoted by $(\Pi_{\lambda},\mathcal{H}_{\lambda})$, is a sub-representation of $(\Pi^{\otimes l},(\mathbb{C}^{n})^{\otimes l})$ for some $l\geq 0$. Thus, we have $(\mathrm{id}_{m}\otimes\Pi_{\lambda})(e_{1})=(\mathrm{id}_{m}\otimes\Pi_{\lambda})(e_{2})$. If $\lambda_{n}<0$, let $\lambda^{\prime}=(\lambda_{1}-\lambda_{n},\dots,\lambda_{n-1}-\lambda_{n},0)$. Then, $(\Pi_{\lambda},\mathcal{H}_{\lambda})$ is equivalent to the tensor product of $(\Pi_{\lambda^{\prime}},\mathcal{H}_{\lambda^{\prime}})$ and $(\Pi_{(-1,\dots,-1)}^{\otimes-\lambda_{n}},\mathcal{H}_{(-1,\dots,-1)}^{\otimes-\lambda_{n}})$, where $\mathcal{H}_{(-1,\dots,-1)}=\mathbb{C}$ and $\Pi_{(-1,\dots,-1)}$ is given as the inverse of the determinant. Since the second tensor components of $e_{1},e_{2}$ are group-like, we have
$(\mathrm{id}_{m}\otimes\Pi_{\lambda})(e_{1})=(\mathrm{id}_{m}\otimes\Pi_{\lambda^{\prime}}\otimes\Pi_{(-1,\dots,-1)}^{\otimes-\lambda_{n}})(e_{1})=(\mathrm{id}_{m}\otimes\Pi_{\lambda^{\prime}}\otimes\Pi_{(-1,\dots,-1)}^{\otimes-\lambda_{n}})(e_{2})=(\mathrm{id}_{m}\otimes\Pi_{\lambda})(e_{2}).$
Therefore, we conclude that $(\mathrm{id}_{m}\otimes\Pi_{\lambda})(e_{1})=(\mathrm{id}_{m}\otimes\Pi_{\lambda})(e_{2})$ for any $\lambda$, and hence $e_{1}=e_{2}$. ∎
Unlike the general setup, we can show the following:
Proposition 6.2.
$K_{m,\rho}$ coincides with $\overline{\Pi_{m,\infty}(B_{m,\rho})}$.
Proof.
By Proposition 5.1, $\Pi_{m,\infty}(B_{m,\rho})$ is contained in $K_{m,\rho}$. Combining Theorem 3.1 and Lemma 6.1, every element in $\mathrm{ex}(K_{m,\rho})$ can be approximated by $\Pi_{m,\infty}(\mathbb{M}_{m}(\mathbb{C})_{+}\otimes\mathcal{E}_{\rho}(U(N)))$. Thus, by Lemma 3.1, $K_{m,\rho}$ is contained in $\overline{\Pi_{m,\infty}(B_{m,\rho})}$. ∎
Let $\rho\in\mathbb{M}_{n}(\mathbb{C})^{*}_{+}$ be faithful. Based on the previous results, a new understanding of separable elements in $(\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C}))_{+}$ is given as follows:
Theorem 6.1.
The set of all separable elements in $(\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C}))_{+}$ is equal to $\overline{\Pi_{m,1}(B_{m,\rho})}$.
Proof.
By Corollary 5.1, $\Pi_{m,1}(B_{m,\rho})$ is contained the set of separable elements. On the other hand, if $x\in(\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C}))_{+}$ is separable, then there exists a $l$-sub-extension $x_{l}$ of $x$ for all $l\geq 2$. Moreover, we may assume that $(x_{l})_{l=0}^{\infty}\in K_{m,\rho}$, where $x_{1}:=x$ and $x_{0}:=(\mathrm{id}_{m}\otimes\rho)(x)$. Thus, by Proposition 6.2, we have $(x_{l})_{l=0}^{\infty}\in\overline{\Pi_{m,\infty}(B_{m,\infty})}$, and hence $x\in\overline{\Pi_{m,1}(B_{m,\rho})}$. ∎
We remark that the quantum de Finetti theorem was not used in the above proof. Thus, we obtain a second proof of Theorem 3.2 based on the results in this section.
Second proof of Theorem 3.2.
We show only that $x\in(\mathbb{M}_{m}(\mathbb{C})\otimes\mathbb{M}_{n}(\mathbb{C}))_{+}$ is separable if there exists a $l$-sub-extension of $x_{l}$ of $x$ for all $l\geq 2$. We may assume that $(x_{l})_{l=0}^{\infty}\in K_{m,\rho}$, where $x_{1}=x$ and $x_{0}:=(\mathrm{id}_{m}\otimes\rho)(x)$. By Proposition 6.2, we have $(x_{l})_{l=0}^{\infty}\in\overline{\Pi_{m,\infty}(B_{m,\infty})}$, and hence $x$ is separable. ∎
Appendix A The infinite dimensional setting
Up to now in this paper, we considered only finite-dimensional $C^{*}$-algebras. In this appendix, we show how to extend our results to general $C^{*}$-algebras. Recall that in Section 3 and 4, we work with the $C^{*}$-algebra $\mathbb{M}_{m}(\mathbb{C})$, or work precisely with the dual $\mathbb{M}_{m}(\mathbb{C})^{*}$, identified to the algebra by the Hilbert–Schmidt inner product.
A.1. A generalized quantum de Finetti theorem for general $C^{*}$-algebras
We first collect the basic concepts in the theory of $C^{*}$-algebras we will need and were not introduced in Section 2. Our main references will be [4, 5, 10].
Let $\mathcal{C}\subset B(\mathcal{H})$ and $\mathcal{D}\subset B(\mathcal{K})$ be two $C^{*}$-algebras acting on the Hilbert spaces $\mathcal{H}$ and $\mathcal{K}$. The algebraic tensor product $\mathcal{C}\odot\mathcal{D}$ of $\mathcal{C}$ and $\mathcal{D}$ is naturally a *-subalgebra of $B(\mathcal{H}\otimes\mathcal{K})$, and its norm-closure, denoted by $\mathcal{C}\otimes\mathcal{D}$, is the minimal tensor product of $\mathcal{C}$ and $\mathcal{D}$.
Let $\mathcal{A}$ and $\mathcal{B}$ be two $C^{*}$-algebras and $l$ be a positive integer. As in Section 3, for any $\sigma\in S_{l}$, we denote also by $\sigma$ the automorphism of $\mathcal{A}\otimes\mathcal{B}^{\otimes l}$, given for any $a\in\mathcal{A}$, $b_{1},\dots,b_{l}\in\mathcal{B}$ by
$\sigma(a\otimes b_{1}\otimes\cdots\otimes b_{l}):=a\otimes b_{\sigma(1)}\otimes\cdots\otimes b_{\sigma(l)}.$
By duality, the action of $S_{k}$ reduces an action on $(\mathcal{A}\otimes B^{\otimes l})^{*}$. We will denote by $(\mathcal{A}\otimes\mathcal{B}^{\otimes l})^{*,S_{l}}$ the subspace of $S_{l}$-invariant elements of $(\mathcal{A}\otimes\mathcal{B}^{\otimes l})^{*}$, and by $(\mathcal{A}\otimes\mathcal{B}^{\otimes l})^{*,S_{l}}_{+}$ the intersection of $(\mathcal{A}\otimes\mathcal{B}^{\otimes l})^{*,S_{l}}$ and $(\mathcal{A}\otimes\mathcal{B}^{\otimes l})^{*}_{+}$.
For any $b\in B_{+}$, let $\mu^{l}_{b}\colon\mathcal{A}\otimes\mathcal{B}^{\otimes l}\to\mathcal{A}\otimes\mathcal{B}^{\otimes(l+1)}$ denote the linear map given by $\mu^{l}_{b}(c)=c\otimes b$ for all $c\in\mathcal{A}\otimes\mathcal{B}^{\otimes l}$. By duality, we get a positive linear map $(\mu^{l}_{b})^{*}\colon(\mathcal{A}\otimes\mathcal{B}^{\otimes(l+1)})^{*}\to(\mathcal{A}\otimes\mathcal{B}^{\otimes l})^{*}$ given by $(\mu^{l}_{b})(\psi):=\psi\circ\mu^{l}_{b}$ for $\psi\in(\mathcal{A}\otimes\mathcal{B}^{\otimes(l+1)})^{*}$. Then, $(\mu^{l}_{b})^{*}((\mathcal{A}\otimes\mathcal{B}^{\otimes(l+1)})^{*,S_{l+1}})\subset(\mathcal{A}\otimes\mathcal{B}^{\otimes l})^{*,S_{l}}$.
Generalizing the setting of Section 3 and keeping the above notations, we defined the following convex cone
$K_{\mathcal{A}^{*},b}:=\left\{(\chi_{l})_{l=0}^{\infty}\in\prod_{l=0}^{\infty}((\mathcal{A}\otimes\mathcal{B}^{\otimes l})^{*})^{S_{l}}_{+}\,\middle|\,(\mathrm{id}_{\mathcal{A}}\otimes\mathrm{id}_{\mathcal{B}^{\otimes l}}\otimes b)(\chi_{l+1})\leq\chi_{l}\text{ for all }l\geq 0\right\}.$
Endowing $\prod_{l=0}^{\infty}(\mathcal{A}\otimes\mathcal{B}^{l})^{*}$ with the topology of component-wise weak-convergence, $K_{\mathcal{A}^{*},b}$ is a closed convex cone.
Lemma A.1.
If $b\in\mathcal{B}_{+}$ is invertible, then $K_{\mathcal{A}^{*},b}$ is the closed convex hull of $\mathrm{ex}(K_{\mathcal{A}^{*},b})$.
Proof.
We only need to prove that for any $r>0$,
$C_{r}:=\{\chi=(\chi_{l})_{l=0}^{\infty}\in K_{\mathcal{A}^{*},b}\mid x_{0}(1_{\mathcal{A}})\leq r\}$
is compact, as the rest of the proof is as in Lemma 3.1. Since $b$ is invertible, there exists $\epsilon>0$ such that $b>\epsilon 1_{\mathcal{B}}$. For every $l\geq 0$, since $\chi_{l}$ is a positive functional, we have
$\|\chi_{l}\|=\chi_{l}(1_{\mathcal{A}}\otimes 1_{\mathcal{B}}^{\otimes l})\leq\epsilon^{-l}\chi_{l}(1_{\mathcal{A}}\otimes b^{\otimes l})\leq\epsilon\chi_{0}(1_{\mathcal{A}})\leq\epsilon^{-l}r.$
By the Banach–Alaoglu theorem, the ball $B_{l,\epsilon^{-l}r}$ of radius $\epsilon^{-l}r$ in $(\mathcal{A}\otimes\mathcal{B}^{\otimes l})^{*}$ are compact. Hence, $\prod_{l\geq 0}B_{l,\epsilon^{-l}r}$ is compact, and $C_{r}\subset\prod_{l\geq 0}B_{l,\epsilon^{-l}r}$ being closed, is also compact. ∎
Using as in Section 3, the results of Chapter 13 of [11], we get the following representation theorem:
Proposition A.1.
Every $\chi\in K_{\mathcal{A}^{*},b}\backslash\{0\}$ can be represented by a maximal Borel probability measure on $\mathrm{ex}(K_{\mathcal{A}^{*},b})$ (in the sense of Pro. 3.1).
By a similar proof as in Theorem 3.1, we have
Theorem A.1.
$\mathrm{ex}(K_{\mathcal{A}^{*},b})$ is contained in $\{(\chi\otimes\beta^{\otimes l})_{l=0}^{\infty}\mid\chi\in\mathcal{A}^{*}_{+},\beta\in\mathcal{B}^{*}_{+}\}$.
We extend Definition 3.1 as follows:
Definition A.1.
Let $\mathcal{A}$ and $\mathcal{B}$ be $C^{*}$-algebras. (1) A positive linear form $\psi\in(\mathcal{A}\otimes\mathcal{B})^{*}_{+}$ is separable if $\psi$ belongs to the closed convex hull of $\mathcal{A}^{*}_{+}\otimes\mathcal{B}^{*}_{+}$. (2) For $b\in\mathcal{B}_{+}$ and any $l\geq 1$, an element $\varphi$ of $(\mathcal{A}\otimes\mathcal{B}^{\otimes l})^{*,S_{l}}_{+}$ is a $l$-sub-extension of $\psi\in(\mathcal{A}\otimes\mathcal{B})^{*}_{+}$ with respect to $b$ if $(\mu^{l}_{b})^{*}(\varphi)\leq\psi$. (3) An element $\psi\in(\mathcal{A}\otimes\mathcal{B})^{*}_{+}$ is $l$-sub-extendable (with respect to $b$) if there is a $l$-sub-extension.
Replacing Lemma 3.1 and Theorem 3.1 by Lemma A.1 and Theorem A.1, the following theorem is proved as Theorem 3.2.
Theorem A.2 (Generalized quantum de Finetti theorem).
Let $\mathcal{A}$ and $\mathcal{B}$ be $C^{*}$-algebras, $b\in\mathcal{B}_{+}$ be invertible and $\psi\in(\mathcal{A}\otimes\mathcal{B})^{*}_{+}$. Then, $\psi$ is separable if and only if $\psi$ is $l$-sub-extendable with respect to $b$ for all $l\geq 1$.
A.2. Infinite-dimensional Biane–Choquet–Deny theorem
The results that we proved in Section 4 can also be extended to a more general $C^{*}$-algebra context.
Let $G$ be a compact group. We keep the notations introduced in Section 4 and fix a positive linear form $\rho\in\mathcal{U}(G)_{*,+}$. For any $m\geq 0$ we defined in Definition 4.2 a positive map $P_{m,\rho}$ on $\mathbb{M}_{m}(\mathbb{C})\otimes\mathcal{U}(G)$.
If $\mathcal{A}$ is a unital $C^{*}$-algebra, we extend Definition 4.2 to a map $P_{\mathcal{A}^{*},\rho}$ defined on the tensor product (as linear spaces) of $\mathcal{A}^{*}\otimes\mathcal{U}(G)$ by $P_{\mathcal{A}^{*},\rho}:=\mathrm{id}_{\mathcal{A}^{*}}\otimes P_{\rho}$.
We say that $x\in\mathcal{A}_{*}\otimes\mathcal{U}(G)$ is positive if $(\mathrm{id}_{\mathcal{A}^{*}}\otimes\pi)(x)$ is a positive linear form in $\mathcal{A}^{*}\otimes B(\mathcal{H})=(\mathcal{A}\otimes B(\mathcal{H}))^{*}$ for any finite-dimensional unitary representation $(\pi,\mathcal{H})$ of $G$. Then, by extending the proof of Proposition 4.1, we can check that $P_{\mathcal{A}^{*},\rho}$ is a positive map.
Extending the definition given in Equation (4.1), we consider the following convex cone
$B_{\mathcal{A}^{*},\rho}:=\{x\in(\mathcal{A}^{*}\otimes\mathcal{U}(G))_{+}\mid P_{\mathcal{A}^{*},\rho}(x)\leq x\}.$
To generalize the statements and the proofs of Lemma 4.1, Proposition 4.2, and Theorem 4.1, recall (see Section 4.1) that if $((\pi_{\alpha},\mathcal{H}_{\alpha}))_{\alpha\in\widehat{G}}$ is the set of irreducible representations of $G$, then each $\pi_{\alpha}$ extends to a unital *-homomorphism from $\mathcal{U}(G)$ to $B(\mathcal{H}_{\alpha})$, and that $\mathcal{U}(G)$ is *-isomorphic to $\prod_{\alpha\in\widehat{G}}B(\mathcal{H}_{\alpha})$ by $\Pi_{G}(x):=(\pi_{\alpha}(x))_{\alpha\in\widehat{G}}$. Therefore, $\Pi_{\mathcal{A}^{*},G}:=\mathrm{id}_{\mathcal{A}^{*}}\otimes\Pi_{G}$ is an linear isomorphism from $\mathcal{A}^{*}\otimes\mathcal{U}(G)$ onto $\mathcal{A}^{*}\otimes\prod_{\alpha\in\widehat{G}}B(\mathcal{H}_{\alpha})\subset\prod_{\alpha\in\widehat{G}}\mathcal{A}^{*}\otimes B(\mathcal{H}_{\alpha})$.
For any $\alpha\in\widehat{G}$, $H_{\alpha}$ being finite dimensional, we identify $\mathcal{A}^{*}\otimes B(\mathcal{H}_{\alpha})$ with $(\mathcal{A}\otimes B(\mathcal{H}_{\alpha}))^{*}$, and we endow $\prod_{\alpha\in\widehat{G}}\mathcal{A}^{*}\otimes B(\mathcal{H}_{\alpha})$ with the product topology of $(\mathcal{A}\otimes B(\mathcal{H}_{\alpha}))^{*}$, endowed with the wak-topology. By considering the topology on $\mathcal{A}^{*}\otimes\mathcal{U}(G)\subset(\mathcal{A}\otimes\mathbb{C}[G])^{*}$ given by the weak topology on $(\mathcal{A}\otimes\mathbb{C}[G])^{*}$, the $\Pi_{\mathcal{A}^{*},G}$ is a topological isomorphism.
Generalizing Lemma 4.1 Proposition 4.2, and their proof, we have
Proposition A.2.
$B_{\mathcal{A}^{*},\rho}$ is the closed convex hull of $\mathrm{ex}(B_{\mathcal{A}^{*},\rho})$. Moreover, every $x\in B_{A^{*},\rho}\backslash\{0\}$ is represented by a maximal Borel probability measure on $\mathrm{ex}(B_{\mathcal{A}^{*},\rho})$ (in the sense of Prop. 3.1).
Proof.
For any $\alpha\in\widehat{G}$ we define a linear functional $1_{A}\otimes\mathrm{Tr}_{\alpha}$ on $\mathcal{A}^{*}\otimes B(\mathcal{H}_{\alpha})$ by
$(1_{\mathcal{A}}\otimes\mathrm{Tr}_{\alpha})(y):=[(\mathrm{id}_{\mathcal{A}}\otimes\mathrm{Tr}_{\alpha})(y)](1_{\mathcal{A}})\quad(y\in\mathcal{A}^{*}\otimes B(\mathcal{H}_{\alpha})),$
where $\mathrm{Tr}_{\alpha}$ is the (non-normalized) trace on $B(\mathcal{H}_{\alpha})$. Moreover, for every $c=(c_{\alpha})_{\alpha\in\widehat{G}}\in\mathbb{R}_{>0}^{\widehat{G}}$ we define $\varphi_{c}\colon(\mathcal{A}^{*}\otimes\mathcal{U}(G))_{+}\to[0,\infty]$ by
$\varphi_{c}(x):=\sum_{\alpha\in\widehat{G}}c_{\alpha}(1_{A}\otimes\mathrm{Tr}_{\alpha})(\Pi_{\mathcal{A}^{*},\alpha}(x))\quad(x\in(\mathcal{A}^{*}\otimes\mathcal{U}(G))).$
To prove the statement, it suffices to show that $C_{\varphi_{c}}:=\{x\in B_{\mathcal{A}^{*},\rho}\mid\varphi_{c}(x)\leq 1\}$ is compact. In fact, the remaining is the same as in Section 4. If $x\in C_{\varphi_{c}}$, then for every $\alpha\in\widehat{G}$, we have
$\|\Pi_{\mathcal{A}^{*},\alpha}(x)\|=|[\Pi_{\mathcal{A}^{*},\alpha}(x)](1_{\mathcal{A}}\otimes 1_{B(\mathcal{H}_{\alpha})})|=|(1_{\mathcal{A}}\otimes\mathrm{Tr}_{\alpha})(\Pi_{A^{*},\alpha}(x))|\leq c_{\alpha}^{-1}.$
Thus, we have $C_{\varphi_{c}}\subset\{x\in\mathcal{A}^{*}\otimes\mathcal{U}(G)\mid\|\Pi_{\mathcal{A}^{*},\alpha}(x)\|\leq c_{\alpha}^{-1}\}$, and hence $C_{\varphi_{c}}$ is compact. ∎
By assuming as in Section 4 that $\rho$ is generating, and by a proof similar to Theorem 4.1, we obtain the following result.
Theorem A.3 ($\mathcal{A}^{*}$-valued Biane–Choquet–Deny theorem).
The extreme points $\mathrm{ex}(B_{\mathcal{A}^{*},\rho})$ are contained in $\mathcal{A}^{*}_{+}\otimes\mathcal{E}_{\rho}(G)$. In particular, $B_{\mathcal{A}^{*},\rho}$ is the closed convex hull of $\mathcal{A}^{*}_{+}\otimes\mathcal{E}_{\rho}(G)$.